OpenStudy (anonymous):

In two years, Linda will be half as old as her mother. Her mother is 38 years old now. Let L = the age of Linda now. Which formula below would equal Linda's current age?

OpenStudy (anonymous):

l+2 = m+2/2 , m=38 , l = 18 .

OpenStudy (anonymous):

where l is lisa, and m is mom.

OpenStudy (anonymous):

yea i got it

OpenStudy (anonymous):

if you want to sum it up, l = (m+2/2)-2

OpenStudy (anonymous):

i think the formula goes like this 2*(l +2) = (m+2)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

lol, thats what i wrote above...

OpenStudy (anonymous):

oh is the same i just dont understood this part (m+2/2) in ur formula it semms to me like (2/2 = 1)

OpenStudy (anonymous):

no, the age of the mother is = 38 , 38+2/2 = 20 not 1.